OpenStudy (anonymous):

Which set of side lengths can form a right triangle? six, seven, eight ten, fourteen, thirty-six twelve, thirty-five, thirty-seven thirteen, twenty-four, thirty

OpenStudy (tkhunny):

Have you considered applying the Pythagorean Theorem? 3, 4, 5??? 3^2 + 4^2 = 9 + 16 = 25 = 5^2 -- YES!!!

OpenStudy (anonymous):

..........wat...?